A plane is flying at 300km/h [S 39 E]. The wind is blowing at 60 km/h [N 75 E]. What is the velocity of the plane relative to the ground?
1
Expert's answer
2019-10-15T06:18:51-0400

Angle between the plane and wind is


"180-75-39=66\degree"

Now relative velocity of plane wrt. ground is


"v_{pg}=v_{pw}+v_{wg}=\sqrt{300^2+60^2+2\times 300\times 60\cos66\degree}"

"329\ km\ h^{-1}"

For direcrion-


"\theta=tan^{-1}(\frac{60sin66\degree}{300+60cos66\degree})=9.584\degree"



So the direction is "9.584\degree" anticlockwise from the direction of plane .